OpenStudy (masterlj2099):

How many solutions will this system of equations have? x − 2y = 24 3x − 6y = 72 No solution Infinite solutions One solution Two solutions

OpenStudy (anonymous):

One second.

OpenStudy (anonymous):

For example, if you put x − 2y = 24 convert that to x = 2y+24, and then put that into your second equation, 3x − 6y = 72. What do you get then @masterlj2099

OpenStudy (trojanpoem):

Infinity

OpenStudy (anonymous):

@TrojanPoem let them figure it out

OpenStudy (anonymous):

@masterlj2099 Can you do the math I just told you to do?

OpenStudy (trojanpoem):

Excuse me, I thought it was easy to figure out that the second equation is simply the first scaled with 3 3(x-2y) = 3 * 24

OpenStudy (masterlj2099):

infinity

OpenStudy (anonymous):

Yeah, but I want @masterlj2099 to do the work out instead of telling him the answer...

OpenStudy (anonymous):

@masterlj2099 can you walk us through how you came up with that answer...sorry to be a pest but policy

OpenStudy (trojanpoem):

@sofia.gulick , You can find it using Gauss elimination, Creamer's rule,....

OpenStudy (trojanpoem):

Or brain.exe

OpenStudy (anonymous):

Okay, you know what I am leaving...because if you don't want my help that is fine.